# DiFSD: Ego-Centric Fully Sparse Paradigm with Uncertainty Denoising and Iterative Refinement for Efficient End-to-End Self-Driving

### v3

@@ 2.1 End-to-End Perception
Besides, Multi-Object Tracking (MOT) across multi-cameras is also required for downstream tasks. Recent works [44, 46, 42, 32, 35] seek to explore the joint detection and tracking methods by introducing track queries to detect the unique instances continuously. Sparse4Dv3 [24] proposes an advanced 3D detector which takes the spatial-temporal information to propagate temporal instances for identity reserving, thus achieving superior end-to-end performance without additional tracking designs.
@@ 3.3 Sparse Perception
Detection. Following the previous sparse perception methods [25, 24], surrounding agents can be represented by a group of instance features and anchor boxes respectively. And each anchor box is denoted as , which contains location, dimension, yaw angle as well as velocity respectively. Taking , and the visual features as input, decoders are adopted to consecutively refine the anchor boxes and update the instance features through deformable aggregation of sample features projected from key points of . The updated instance features are adopted to predict the classification scores and box offsets respectively. Temporal instance denoising is introduced to improve model stability.
@@ 3.6 Uncertainty Denoising
Due to the planning-oriented modular design, output uncertainty from each individual module will be inevitably introduced and passed through to the downstream tasks, leading to inferior and fragile system. Under this circumstance, we propose a two-level uncertainty modeling strategy to further stabilize the whole framework. On one hand, position-level diffusion process is performed on ground-truth boxes of interactive agents for additional trajectory prediction of noisy agents equipped with groups of random noises following uniform distributions. locates within two different ranges of and to indicate positives and negatives respectively, where indicates the noise scale. This process aims to promote the stability of motion forecasting for interactive agents with uncertain detected positions, scales and velocities. On the other hand, trajectory-level denoising process is also introduced for robust offset prediction of proposal trajectory of ego-vehicle in the planning optimization stage. Different from the position diffusion of detection or motion query described above, we apply the random noise to trajectory offsets of ego-vehicle , where depends on the Final Displacement (FD) of ground-truth ego future trajectory.
3.7 End-to-End Learning
4 Experiments
4.1 Datasets and Metrics
4.2 Implementation Details

### v4

@@ 3.3 Sparse Perception
Detection. Following the previous methods [25, 24], surrounding agents can be represented by a group of instance features and anchor boxes respectively. And each anchor box can be denoted as:
which contains location, dimension, yaw angle as well as velocity respectively. Taking , and the visual features as input, decoders are adopted to consecutively refine the anchor boxes and update the instance features through deformable aggregation of sample features projected from key points of . The updated instance features are adopted to predict the classification scores and box offsets respectively. Temporal instance denoising is introduced to improve model stability.
@@ 3.6 Uncertainty Denoising
Due to the planning-oriented modular design, output uncertainty from each individual module will be inevitably introduced and passed through to the downstream tasks, leading to inferior and fragile system. Under this circumstance, we propose a two-level uncertainty modeling strategy to further stabilize the whole framework.
On one hand, position-level diffusion process is performed on ground-truth boxes of interactive agents for additional trajectory prediction of noisy agents:
which are equipped with groups of random noises following uniform distributions. locates within two different ranges of and to indicate positives and negatives respectively, where indicates the noise scale. This process aims to promote the stability of motion forecasting for interactive agents with uncertain detected positions, scales and velocities.
On the other hand, trajectory-level denoising process is also introduced for robust offset prediction of proposal trajectory of ego-vehicle in the planning optimization stage. Different from the position diffusion of detection or motion query described above, we apply the random noise to trajectory offsets of ego-vehicle , where depends on the Final Displacement (FD) of ground-truth ego future trajectory.
@@
3.7 End-to-End Learning
4 Experiments
4.1 Datasets and Metrics
4.2 Implementation Details